# Establishing a Foundation for Tetun Text Ad-Hoc Retrieval: Stemming, Indexing, Retrieval, and Ranking

### v2

@@ 6.5. Discussions
In higher cutoffs, network-based methods still outperformed traditional term weighting approaches, with in-degree consistently delivering the best results for Tetun and Portuguese, showing improvements of up to +0.18 points. However, in English, the degree marginally surpassed in-degree. These results indicate that network-based methods maintain a stronger advantage as the stopword list expands, except in English at P@1000, where all methods produced identical scores, likely due to characteristics of the dataset.
@@
Overall, in Tetun, in-degree proved slightly more effective than degree, while in English, degree marginally outperformed in-degree. In Portuguese, both methods performed similarly (see Figure 2). Since degree is the sum of in-degree and out-degree, and its performance is comparable to in-degree, in-degree offers computational efficiency advantages.
@@
Figure 2. Comparison of the Network-Based Approach Performance at Mid-Range and Higher Cutoff Levels.
[…]
Comparison of the Network-Based Approach Performance at Mid-Range and Higher Cutoff Levels.
@@
For traditional term-weighting approaches, IDF outperformed TF and TF-IDF at higher cutoffs in Tetun. At mid-range cutoffs, IDF and TF achieved identical scores, while at lower cutoffs, TF, IDF, and TF-IDF all produced identical scores. In contrast, for Portuguese and English, TF, IDF, and TF-IDF yielded identical results across all cutoffs. This consistency in Portuguese and English may be due to the more structured and mature linguistic resources available for these languages, such as well-established stopword lists and corpora, which minimize variations in term weighting effectiveness. In Tetun, the language’s lesser-resourced nature likely results in greater sensitivity to different weighting methods, leading to performance differences at higher cutoffs.
Furthermore, the total number of stopwords we developed for Tetun is comparable to other LRLs such as Marathi (99 stopwords), Bengali (114 stopwords), and Hindi (165 stopwords) (Sahu et al., 2023); and Kinyarwanda (80 stopwords) and Kirundi (59 stopwords) (Niyongabo et al., 2020).
6.6. Conclusions
@@
This study highlights the superiority of network-based approaches, particularly in-degree and degree, over traditional term weighting methods for identifying stopwords in both high- and low-resource languages, especially when dealing with larger stopword sets. Although traditional term weighting and network-based methods perform comparably at smaller cutoffs (up to 25 terms), network-based approaches demonstrate greater effectiveness as the number of evaluated terms increases. For smaller stopword lists, the differences between methods are less significant. However, when working with lists of 25 or more stopwords, network-based approaches are recommended for their superior performance at mid-range and higher cutoffs. The consistent in-degree performance observed in Tetun is aligned with the findings reported by Ali et al. (Ali et al., 2024), further validating the effectiveness of network-based methods for stopword detection tasks, specifically in LRL scenarios.
@@ 7. Labadain-Stemmer Development
This section outlines the development of Labadain Stemmer, a stemming algorithm specifically designed for Tetun. It covers the identification of Tetun affixes and the development of Labadain Stemmer variants tailored to the language. Additionally, it details the process of generating sample words, which were evaluated by humans to serve as ground truth for assessing the accuracy of the Labadain Stemmer. The section further presents the experiments conducted, their results, and corresponding discussions, concluding with a summary of limitations and key observations.
@@ 7.1. Tetun Affixes
This study focuses on commonly used affixes in Tetun (van Klinken et al., 2002; Greksáková, 2018; Hull and Correia, 2005), excluding circumfixes and reduplication due to their limited productivity and usage, as discussed in subsection 3.3. The native Tetun prefixes are “ha”“nak”, and “nam”, while the native suffixes comprises “n”“-nain”“-teen”, and “dór”. Additionally, Portuguese-derived suffixes are adapted from the list of Portuguese suffixes used in the Portuguese stemmer in Snowball (Snowball, 2005a), as presented in Table 29 in the Appendix 12.2. Since Tetun has few inflectional affixes, stemming native Tetun words is a straightforward process that involves matching words with a predefined list of Tetun prefixes and suffixes at the beginning and end.
@@ 7.2. Stemmer Variants
Tetun consists of both native words and a significant number of Portuguese loanwords, particularly verbs, nouns, and adjectives (Hajek and van Klinken, 2019; Greksáková, 2018; van Klinken and Hajek, 2018; van Klinken et al., 2002). To address this linguistic mix, the Tetun stemmer is designed with three variants: lightmoderate, and heavy. Each variant is detailed in the following subsections.
@@ 7.2.1. Light Stemmer
The light stemmer is specifically designed to reduce the suffixes of Portuguese-derived words used in Tetun to their root forms. This variant adapts the Portuguese stemmer from Snowball, using a customized list of Portuguese-derived suffixes, transformed according to the loanword transformation rules established by the INL (of Linguistics, INL)Table 10 outlines some rules for transforming Portuguese-derived words according to the INL standardization.
@@
The Tetun light stemmer is a simplified version of the Portuguese stemmer, adapted to handle loanwords in a way that aligns with Tetun’s unique morphological characteristics. In the light stemmer, the algorithm maintains the linguistic regions used in the Portuguese stemmer, which were derived from the Spanish stemmer in Snowball (Snowball, 2005b). The definitions of the linguistic regions used in the Tetun light stemmer algorithm are provided in Table 28 in the Appendix 12.2.
Tetun light stemmer processes words sequentially using suffix lists developed according to the characteristics of Portuguese loanwords in Tetun (see Table 29). The stemming procedure is summarized below, with the full algorithm detailed in Algorithm 12.2:
@@
    Word length validation: After receiving an input word, the algorithm begins by validating its length. If the word has fewer than four characters, it is returned without stemming.
@@
    Standard suffix removal: If the word is longer than three characters, the algorithm searches for the longest matching suffix from the general suffix list. If a matching suffix is found in a specific region of the word, it is deleted or replaced.
@@
    Verb suffix removal: If no suffix is removed in step (2), the algorithm checks a list of verb-specific suffixes and removes them if appear in the appropriate region of the word.
@@
    Residual suffix removal: If neither of the previous steps results in suffix removal, the algorithm looks at the remaining simple suffixes list and removes it as a final step.
@@
    Return original word: If none of the steps results in suffix removal, the input word is returned unchanged.
@@ 7.2.2. Moderate Stemmer
The moderate stemmer builds on the light stemmer variant by removing suffixes from both Portuguese loanwords and native Tetun words. It follows the same algorithm as the light stemmer (detailed in Algorithm 12.2) with an additional step, labeled (3.1), to process native Tetun suffixes, inserted between steps (3) and (4).
@@ 7.2.3. Heavy Stemmer
The heavy stemmer further extends the moderate stemmer variant by also removing native Tetun prefixes. In this variant, the processing of native Tetun prefixes occurs between steps (3.1) and (4).
7.3. Sample Word Generation
This subsection outlines the creation of sample words that are used for intrinsic experimental and evaluation purposes. Selecting sample words to assess stemming performance poses challenges due to potential bias and limited generalization. To mitigate this, we designed a systematic methodology to construct this sample from a dataset containing diverse collection of categories and sources, with human involvement in the loop.
Figure 3. Process of Constructing a Text Sample for Evaluating Tetun Stemmer’s Performance.
[…]
Process of Constructing a Collection for Evaluating Tetun Stemmer.
The process of constructing sample words used for the experiment and evaluation is illustrated in Figure 3. First, we preprocessed the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), which involved lowercase, normalizing apostrophes, removing punctuation, special characters, numbers, and extra spaces. After this, the text was tokenized into individual words using the Tetun tokenizer (de Jesus and Nunes, 2024c), and deduplication was performed to remove duplicate tokens. We then normalized and excluded stopwords, along with tokens shorter than four characters. The Tetun LID model (de Jesus and Nunes, 2024c) was applied with a threshold score of 0.95 to filter candidate sample words and further refine the sample. Finally, the candidate sample words were verified by native Tetun speakers to produce the final sample words.
The sample word verification with human involvement was carried out by six native Tetun-speaking students, consisting of five undergraduate students and one Ph.D. student. Each student was responsible for verifying approximately 2,732 words, sorted in ascending alphabetical order. They checked the correctness of each word using both the INL dictionary (Correia et al., 2005) and the Portuguese loanword dictionary (Greksáková, 2018) as reference materials. In the verification process, a significant number of misspelled words were identified, such as the word “konsiderasaun” (consideration, in English) appearing as [ konsidersaunkonsiderasaunnkonsideransaun ]. Furthermore, some words that originate from Tetun Terik (a dialect of Timor-Leste), Tetun DIT, or others not found in the dictionaries were removed from the final sample words. A summary of the sample words resulting from this process is provided in Table 11.
@@ 7.5. Intrinsic Evaluation
In intrinsic evaluation, the Paice metric (Paice, 1994) was used to evaluate the quality of stemmer variants by measuring how effectively they reduce various word forms to a common root. This metric balances the effects of understemming and overstemming, both of which impact precision and recall in text-processing tasks. In IR, a high understemming can lower recall, resulting in relevant documents not being retrieved, while a high overstemming can decrease precision by retrieving many irrelevant documents.
@@ 7.6. Experimental Setting
To compute the Paice metric, a list of words is organized into conceptual groups based on semantic and morphological relationships. These groups represent the target, and an ideal stemmer should conflate words with these conceptual groupings. The stemmers were applied to the word list, and their performance was evaluated by measuring how well they matched the predefined conceptual groups. Examples of the conceptual grouping are shown below:
@@ 7.7. Evaluation and Result
Figure 4. UI vs. OI Plot Showing ERRT Distances.
[…]
UI vs. OI Plot Showing ERRT Distances.
@@
The UI, OI, SW, and ERRT values are presented in Table 14. As expected, the light variant exhibits the highest understemming value (by +0.007 points), while the moderate and heavy variants have identical lowest ERRT values (both lower by approximately -0.009 points), and all variants have the same overstemming values. Further investigation revealed that the difference in words stemmed by the heavy variant compared to the moderate variant was limited to only six words. This is due to the small proportion of native Tetun words, which makes up only 18.21% of the total (see Table 13), and the limited number of Tetun prefixes (outlined in subsection 7.1).
Regarding the overstemming value, since native Tetun has few inflectional forms in the word list, applying both the moderate and heavy variants had no impact on the overstemming value. The minimal difference in the ERRT values presented in Table 14, indicates that the lightmoderate, and heavy stemmer variants perform quite similarly.
7.8. Discussions
Tetun, as a language with relatively few inflectional affixes (Greksáková, 2018; Hull and Correia, 2005; van Klinken et al., 2002), often includes short affixes, such as the prefix “ha” and the suffix “n”, pose challenges to stemming algorithms in correctly handling native Tetun affixes. Some verbs and nouns begin or end with these characters, though they are not affixes, as seen in words like “halimar” (play), “hariis” (bathe), “aman (father), “inan (mother), “ibun (mouth), “liman (hand), “ulun (head), among others. Additionally, removing these characters from certain words changes their meaning. For example, removing “ha” from “halimar” (play) results in “limar” (rasp), and removing the suffix “n” from “liman (hand) becomes “lima” (five).
Given that three variants of the Labadain-Stemmer show similar performance, it might be affected by the characteristics of the sample words used for evaluation. Factors such as the proportion of native Tetun words, the presence of affixes, and the term distribution within the sample could influence the stemming algorithms’ outcomes. Developing a more balanced sample word composition could provide deeper insight into the observed results.
@@ 7.9. Limitations
The Instituto Nacional de Linguística (INL) launched the Kursu Gramátika Tetun (Tetun Grammar Course) in 2005, which served as a reference for teachers, translators, journalists, and students (Hull and Correia, 2005). It includes several Tetun affixes, such as prefixes [ hak, na, ma ] and suffixes [ -laek, k ]. However, more recent research by Greksáková in 2018 (Greksáková, 2018) reported that many of these affixes have been largely replaced by words such as “sai” (become) and “laiha” (without), making these affixes less productive in Tetun. As the INL has not updated its 2005 publication on Tetun grammar, the current state of Tetun morphology remains unclear.
@@
7.10. Conclusions
This study developed and assessed the effectiveness of the Labadain-Stemmer incorporating suffixes of Portuguese loanwords and the affixes of native Tetun words. The Tetun affixes used were based on those commonly reported by Klinken et al. (van Klinken et al., 2002), the INL (Hull and Correia, 2005), and Greksáková (Greksáková, 2018). To evaluate stemmer performance, we systematically constructed sample words and established the baseline for Labadain-Stemmer, testing three variants (light, moderate, and heavy). Results showed that integrating native Tetun affixes into the stemming process was marginally more effective than focusing solely on the suffixes of Portuguese loanwords.
However, the unbalanced representation of Portuguese loanwords and native Tetun words in the sample set poses a limitation. Future research should address this by using more balanced datasets that adequately represent both Portuguese loanwords and native Tetun words. Furthermore, the involvement of expert linguists specializing in Tetun will be crucial for enhancing the accuracy and overall effectiveness of the Labadain-Stemmer.

### v6

@@ 6.5. Discussion
In higher cutoffs, network-based methods still outperformed traditional term weighting approaches, with in-degree consistently delivering the best results for Tetun and Portuguese, showing improvements of up to +0.18 points. However, in English, the degree marginally surpassed the in-degree. These results indicate that network-based methods maintain a stronger advantage as the stopword list expands, except in English at P@1000, where all methods produced identical scores, likely due to characteristics of the dataset.
@@
Overall, in Tetun, in-degree was slightly more effective than degree, while in English, degree marginally outperformed in-degree. In Portuguese, both methods performed similarly (see Figure 2). Since the degree is defined as the sum of in-degree and out-degree and yields performance comparable to in-degree, the latter offers advantages in computational efficiency.
@@
Figure 2. Comparison of the Network-Based Approach Performance at Mid-Range and Higher Cutoff Levels.:
@@
For traditional term-weighting approaches, IDF outperformed TF and TF-IDF at higher cutoffs in Tetun. At mid-range cutoffs, IDF and TF achieved identical scores, while at lower cutoffs, TF, IDF, and TF-IDF all produced identical scores. In contrast, for Portuguese and English, TF, IDF, and TF-IDF yielded identical results across all cutoffs. This consistency in Portuguese and English may be due to the more structured and mature linguistic resources available for these languages, such as well-established stopword lists and corpora, which minimize variations in term weighting effectiveness. In Tetun, the language’s lesser-resourced nature likely results in greater sensitivity to different weighting methods, leading to performance differences at higher cutoffs. Furthermore, the total number of stopwords we developed for Tetun is comparable to other LRLs such as Marathi (99 stopwords), Bengali (114 stopwords), and Hindi (165 stopwords) (Sahu et al., 2023); and Kinyarwanda (80 stopwords) and Kirundi (59 stopwords) (Niyongabo et al., 2020).
6.6. Conclusion
@@
This study highlights the superiority of network-based approaches, particularly in-degree and degree, over traditional term weighting methods for stopword detection in both high- and low-resource languages, especially when dealing with larger stopword sets. Although traditional term weighting and network-based methods perform comparably at smaller cutoffs (up to 25 terms), network-based approaches demonstrate greater effectiveness as the number of evaluated terms increases. For smaller stopword lists, the differences between methods are less significant. However, when working with lists of 25 or more stopwords, network-based approaches are recommended for their superior performance at mid-range and higher cutoffs. The consistent in-degree performance observed in Tetun is aligned with the findings reported by Ali et al. (2024), further validating the effectiveness of network-based methods for stopword detection tasks, specifically in under-resourced scenarios.
@@ 7. Labadain-Stemmer Development
This section describes the development of Labadain-Stemmer, a stemming algorithm specifically designed for Tetun. It covers the identification of Tetun affixes and the creation of Labadain-Stemmer variants tailored to the language. Additionally, it details the process of generating a sample of words, which native Tetun speakers assessed to serve as the ground truth for evaluating the accuracy of Labadain-Stemmer. Finally, the section presents the experiments conducted, their results, and corresponding discussion, concluding with a summary of limitations and key observations.
@@ 7.1. Tetun Affixes
This study focuses on commonly used affixes in Tetun (van Klinken et al., 2002; Greksáková, 2018; Hull and Correia, 2005), excluding circumfixes and reduplication due to their limited productivity and usage, as discussed in Subsection 3.3. The native Tetun prefixes are “ha”“nak”, and “nam”, while the native suffixes comprises “n”“-nain”“-teen”, and “dór”. Additionally, Portuguese-derived suffixes are adapted from the list of Portuguese suffixes used in the Portuguese stemmer in Snowball (Snowball, 2005a), as presented in Table 30 in the Appendix 12.2. Since Tetun has few inflectional affixes, stemming native Tetun words is a straightforward process that involves matching words with a predefined list of Tetun prefixes and suffixes at the beginning and end of each word.
@@ 7.2. Stemmer Variants
Tetun consists of both native words and a significant number of Portuguese loanwords, particularly verbs, nouns, and adjectives (Hajek and van Klinken, 2019; Greksáková, 2018; van Klinken and Hajek, 2018; van Klinken et al., 2002). To address this linguistic mix, the Labadain-Stemmer is designed with three variants: lightmoderate, and heavy. Each variant is detailed in the following subsections.
@@ 7.2.1. Light Stemmer
The light stemmer is designed to remove suffixes from Portuguese-derived words used in Tetun. This variant adapts the Portuguese stemmer from Snowball, incorporating a customized list of Portuguese suffixes. These suffixes were modified based on the loanword transformation rules defined by the INL (of Linguistics, INL), as detailed in Table 10.
@@
The Tetun light stemmer is a simplified adaptation of the Portuguese stemmer, designed to handle loanwords while accounting for Tetun’s unique morphological characteristics. It retains the linguistic regions utilized in the original Portuguese stemmer, which were adapted from the Spanish stemmer in Snowball (Snowball, 2005b). The definitions of these linguistic regions, as applied in the Tetun light stemmer algorithm, are provided in Table 29 of Subsection 12.2.
The Tetun light stemmer processes words sequentially using a list of suffixes developed to account for the specific features of Portuguese loanwords in Tetun (see Table 30). The stemming procedure is summarized below, with the corresponding algorithm provided in Algorithm 12.2:
@@
    Word length validation: After receiving an input word, the algorithm begins by validating its length. If the word contains fewer than four characters, it is returned without stemming.
@@
    Standard suffix removal: For words longer than three characters, the algorithm searches for the longest matching suffix from the general suffix list. If a matching suffix is found in a specific region of the word, the suffix is deleted or replaced accordingly.
@@
    Verb suffix removal: If no suffix is removed in step (2), the algorithm checks for verb-specific suffixes. It is removed if a matching suffix is found within the appropriate region of the word.
@@
    Residual suffix removal: If neither of the previous steps results in suffix removal, the algorithm looks at the remaining simple suffixes list and removes it as the final step.
@@
    Return original word: If none of the steps result in suffix removal, the input word is returned unchanged.
@@ 7.2.2. Moderate Stemmer
The moderate stemmer extends the functionality of the light stemmer by handling suffixes from both Portuguese loanwords and native Tetun. It adheres to the same algorithm as the light stemmer (outlined in Algorithm 12.2), with the addition of a new step of 4.1, specifically designed to process native Tetun suffixes. This additional step is executed between steps 4 and 5 of the algorithm.
@@ 7.2.3. Heavy Stemmer
The heavy stemmer builds on the functionality of the moderate stemmer by introducing the removal of native Tetun prefixes. In this variant, the processing of native Tetun prefixes is integrated between steps 4.1 and 5 of the algorithm.
7.3. Text Sample Construction for Evaluation
This subsection outlines the creation of sample words that are used for intrinsic experimental and evaluation purposes. Selecting sample words to assess stemming performance poses challenges due to potential bias and limited generalization. To mitigate this, we designed a systematic methodology to construct this sample from a dataset containing a diverse collection of categories and sources, with human involvement in the loop.
Figure 3. Process of Constructing a Text Sample for Evaluating Tetun Stemmer’s Performance.:
The process of constructing sample words used for the experiment and evaluation is illustrated in Figure 3. First, we preprocessed the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), which involved lowercase, normalizing apostrophes, and removing punctuation, special characters, numbers, and extra spaces. After this, the text was tokenized into individual words using the Tetun tokenizer (de Jesus and Nunes, 2024e), and deduplication was performed to remove duplicate words. This preprocessed text was then tokenized into individual words using the Tetun Word Tokenizer (de Jesus and Nunes, 2024d), and deduplication was performed to remove duplicate tokens. Stopwords were normalized and subsequently excluded, along with tokens shorter than four characters. To further refine the candidate sample words, the Tetun LID model (de Jesus and Nunes, 2024c) was applied with a threshold score of 0.95 to filter out words that did not meet this criterion. Finally, the refined candidate sample words were validated by native Tetun speakers to produce the final set of sample words.
The sample word verification with human involvement was carried out by six native Tetun speakers, consisting of one Ph.D. and five undergraduate students. Each student was tasked with verifying approximately 2,732 words, sorted in ascending alphabetical order. They checked the correctness of each word using both the INL dictionary (Correia et al., 2005) and the Portuguese loanword dictionary (Greksáková, 2018) as reference materials. During the verification process, a considerable number of misspelled words were identified, such as the word “konsiderasaun” (consideration, in English) appearing as [ konsidersaunkonsiderasaunnkonsideransaun ]. Additionally, some words originating from Tetun Terik, Tetun DIT, or other variants not present in the reference dictionaries were excluded from the final sample. A summary of the resulting sample of words from this process is presented in Table 11.
@@ 7.5. Intrinsic Evaluation
In intrinsic evaluation, the Paice metric (Paice, 1994) was used to evaluate the stemmer variant quality by measuring how effectively they reduce various word forms to a common root. This metric balances understemming and overstemming effects, both of which impact precision and recall in text-processing tasks. In IR, a high understemming lowers recall, resulting in relevant documents not being retrieved, while a high overstemming hurts precision by retrieving many irrelevant documents.
@@ 7.6. Experimental Setting
To compute the Paice metric, a list of words is first organized into conceptual groups based on semantic and morphological relationships. These groups serve as the target, and an ideal stemmer should conflate words according to these conceptual groupings. The stemmers were then applied to the word list, and their performance was evaluated by measuring how accurately they matched the predefined conceptual groups. Examples of these conceptual groupings are provided below, where the root word is shown on the left side and its corresponding conflated words are listed on the right side, separated by a colon delimiter.
@@ 7.7. Evaluation and Results
Figure 4. UI vs. OI Plot Showing ERRT Distances.:
The UI, OI, SW, and ERRT values are presented in Table 14. As expected, the light variant exhibits the highest understemming value (by +0.007 points), while the moderate and heavy variants have identical lowest ERRT values (both lower by approximately -0.009 points), and all variants have the same overstemming values. Further investigation revealed that the difference in words stemmed from the heavy variant compared to the moderate variant was limited to only six words. This is due to the small proportion of native Tetun words, which make up only 18.21% of the total (see Table 13), and the limited number of Tetun prefixes (outlined in Subsection 7.1).
@@
Regarding the overstemming value, since native Tetun has few inflectional forms in the word list, applying both the moderate and heavy variants had no impact on the overstemming value. The minimal difference in the ERRT values presented in Table 14 indicates that the lightmoderate, and heavy stemmer variants perform quite similarly.
7.8. Discussion
Tetun, as a language with relatively few inflectional affixes (Greksáková, 2018; Hull and Correia, 2005; van Klinken et al., 2002), often includes short affixes, such as the prefix “ha” and the suffix “n”, which present challenges to stemming algorithms in correctly handling native Tetun affixes. Some verbs and nouns begin or end with these characters, though they are not affixes, as seen in words like “halimar” (play), “hariis” (bathe), “aman (father), “inan (mother), “ibun (mouth), “liman (hand), “ulun (head), among others. Additionally, removing these characters from certain words changes their meaning. For example, removing “ha” from “halimar” (play) results in “limar” (rasp), and removing the suffix “n” from “liman (hand) becomes “lima” (five).
Given that three variants of the Labadain-Stemmer show similar performance, it might be affected by the characteristics of the sample of words used for evaluation. Factors such as the proportion of native Tetun words, the presence of affixes, and the term distribution within the sample could influence the stemming algorithms’ outcomes. Developing a more balanced sample of word composition could provide deeper insight into the observed results.
@@ 7.9. Limitations
The Instituto Nacional de Linguística (INL) launched the Kursu Gramátika Tetun (Tetun Grammar Course) in 2005, which served as a reference for teachers, translators, journalists, and students (Hull and Correia, 2005). It includes several Tetun affixes, such as prefixes [ hak, na, ma ] and suffixes [ -laek, k ]. However, more recent research by  Greksáková (2018) in 2018 reported that many of these affixes have been largely replaced by words such as “sai” (meaning “become”) and “laiha” (meaning “without”), making these affixes less productive in Tetun. As the INL has not updated its 2005 publication on Tetun grammar, the current state of Tetun morphology remains unclear.
@@
7.10. Conclusion
This study developed and assessed the effectiveness of the Labadain-Stemmer, incorporating suffixes of Portuguese loanwords and the affixes of native Tetun words. The Tetun affixes used were based on those commonly reported by Klinken et al. (van Klinken et al., 2002), the INL (Hull and Correia, 2005), and Greksáková (2018). To evaluate stemmer performance, we systematically constructed sample words and established a baseline for Labadain-Stemmer, testing three variants (light, moderate, and heavy). Results showed that integrating native Tetun affixes into the stemming process was marginally more effective than focusing solely on the suffixes of Portuguese loanwords.
However, one of the limitations in this study is the unbalanced representation of Portuguese loanwords and native Tetun words in the sample set. Future research should address this by using more balanced datasets that adequately represent both Portuguese loanwords and native Tetun words. Furthermore, the involvement of expert linguists specializing in Tetun will be crucial for enhancing the accuracy and overall effectiveness of the Labadain-Stemmer. To enable reproducibility, the stemmer algorithms have been released under the MIT License (de Jesus and Nunes, 2025b), to encourage further research and development in Tetun information retrieval.